HUBBLE SPACE TELESCOPE DAILY REPORT #2090 PERIOD COVERED: 0000Z (UTC) 03/27/98 - 0000Z (UTC) 03/30/98 Daily Status Report as of 089/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The first iteration of this proposal was not executed due to the STIS suspend. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.2 Completed Five Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Three Sets of STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The first iteration of this proposal was not executed due to the STIS suspend. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed Four Sets of WF/PC-2 6540 (Imaging Circumstellar Nebulae Around Luminous Blue Variables in the Magellanic Clouds) The WF/PC-2 was used to make observations of the curcumstellar nebula around the luminous blue variable stars BE381, HD269858, HD269927C, S61 and S119 which are located in the Magallanic Clouds. The objectives are to study the distribution of the ionized gas and dust in order to determine the morphology and physical properties of the circumstellar nebulae, by observing them in a narrow, line-free V continuum filter and in the H Alpha emission line. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Five Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The first three iterations of this proposal was not executed due to the STIS suspend. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.6 Completed Ten Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.7 Completed WF/PC-2 7385 (Low-Mass Halo Binaries) The WF/PC-2 was used to make observations of the low mass, halo, binary, {Population II}, M-type, sub-dwarf, star LHS407. This star is too faint to be resolved through ground-based speckle interferometry, thus HST represents the only means of probing separations. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Two Sets of NIC/1/2/3 7819 (A New Stellar Generation in 30 Doradus) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the 30-Doradus nebula. There is ample evidence for current star formation within this nebula, which appears to have been triggered by the massive central cluster {R136}, in and just beyond the curved nebular filaments surrounding it. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Eight Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.10 Completed STIS/CCD 7438 (Spatially-Resolved Spectroscopy of the UVX in Elliptical Galaxies) The Space Telescope Imaging Spectrograph (CCD) was to be used to make spatially resolved far-ultraviolet observations of the large elliptical galaxy NGC-3115. The prominent far-UV upturn {UVX} in the spectra of elliptical galaxies originates from stars in their old populations which have evolved beyond the He-flash at the tip of the red giant branch. This proposal was not executed due to the STIS suspend. 1.11 Completed NIC/3 7818 (Spectrophotometry of Pluto and Charon) The Near Infrared Camera (Camera Three) was used to make high signal-to-noise ratio specrophotometric observations of the planet Pluto and its moon Charon. These observations will be used to separate the spectroscopic signatures of each object in future ground-based combined light measurements. These spectra will also provide a very sensitive search for trace molecular components on the surfaces of Pluto and Charon. The observations were completed as scheduled, and no problems were reported. 1.12 Completed STIS/CCD 7810 (Sensitivity of Important Secondary Wavelength Settings) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard source BD+75D325 as part of a calibration for sensitivity of important secondary wavelength settings. This program is the basic sensitivity measurement for four secondary wavelength settings which are heavily used in Cycle 7. It is run once a year. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star 04916+175 as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6770 (Resolving the Nature of Nova Shells) The WF/PC-2 was used to make images of the post nova remnant DK-Lac in order to study its surrounding shells structure. Past data has indicate a strong correlation between nova speed class and the ellipticity of the resulting remnants. As described in HSTAR 6413 and in 2.1, the acquisition for this proposal defaulted to fine lock on one FGS only, possibly affecting three observations. Otherwise, the observations were completed as scheduled, and no problems were noted. 1.15 Completed NIC/1/2 7365 (Unscrambling the Near-Infrared Properties of Young Planetary Nebulae) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young planetary nebula OH231.8+4. Objects of this class display a number of enigmatic features that have prevented a full understanding of post-asymptotic giant branch stellar evolution. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Two Sets of STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed Two Sets of WF/PC-2/STIS/CCD/NIC/1/2 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One and Two) were used to make observations of the rich star cluster NGC-2209 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.18 Completed STIS/CCD/MA1/MA2 7509 (High-Velocity Gas and Dust Evolution in Chamaeleon Clouds) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the star HD102065 in order to study the high-velocity gas and dust evolution in Chamaeleon Clouds. These data should yield an absorption spectra obtained from the nearby star located behind this complex of molecular cloud. A wave length calibration observation was also taken. The observations were completed as planned, and no problems were reported. 1.19 Completed NIC/3 7499 (NICMOS Extragalactic Parallel Grism Survey) The Near Infrared Camera (Camera Three) was used to make observations as part of an extragalactic parallel Grism Survey. This proposal is a parallel program using NICMOS's unique capabilities for slitless spectroscopy in the near IR. The observations were completed as planned, and no anomalies were reported. 1.20 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.21 Completed FGS 7492 (The Distances to Dwarf Novae, and the Calibration of the Technique of Infrared Spectroscopic Parallax) The Fine Guidance Sensors was used to make astrometric observations of the dwarf nova SS-Aug and several other reference stars. The distances to dwarf novae are poorly known. The distance to the object can be measured using a technique of infrared spectroscopic parallax. The observations were completed as scheduled, and no anomalies were noted. 1.22 Completed NIC/1/2/3 7859 (Blue Compact Dwarf Galaxies -- Young or Old?) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the compact young dwarf galaxy UGC-5272. These blue compact dwarf galaxies challenge models of galaxy formation to explain the presence of small, low-metallicity galaxies experiencing vigorous bursts of star formation. The observations were completed as scheduled, and no problems were reported. 1.23 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-3623. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.24 Completed NIC/1/2 7454 (The Stellar Content of Powerful Radio Galaxies -- A NICMOS Survey of 0.8 < z < 1.8 3CR Sources) The Near Infrared Camera (Cameras One and Two) was used to make survey observations of the powerful radio galaxy 3C267. The study of distant radio galaxies is complicated by their powerful active nucleus, whose emission dominates at rest-frame ultraviolet wavelengths, producing the observed peculiar, elongated morphologies; the underlying stellar component is lost in the glare. The observations were completed as planned, and no anomalies were reported. 1.25 Completed STIS/MA2 7720 (MAMA Full Field Sensitivity Monitor and PSF Check) The Space Telescope Imaging Spectrograph (MA2) was used to observations of object NGC-6681 to monitor the sensitivity of the FUV-MAMA and NUV-MAMA over the full field and monitor the PSF. The observations completed as planned, with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 23 Successful : 23 The acquisition at 087/031317Z defaulted to fine lock backup on FGS-1 only. HSTAR 6413 was written. The proposal mentioned in 1.14 may have been affected. Scheduled Reacquisitions : 20 Successful : 20 2.2 FHST Updates: Scheduled : 47 Successful : 47 2.3 Operations Notes: TTAC1 went down at 086/0957Z. The ground system was back on-line at 086/1230Z with TTAC2. Using ROP NS-3, the NSSC-1 status buffer was dumped at 086/1415Z and at 086/1442Z. The STIS (and MAMA1) was recovered successfully from suspend at 086/1723Z via operations request #13656. MAMA2 recovery was accomplished via the SMS at 088/013805Z. The proposals described in 1.1, 1.3, 1.5, and 1.10 were affected by this event. Analysis of the events surrounding the suspend indicates that the lamp illumination was not bright enough to trigger the hardware bright scene detected (BSD) monitor. The global event threshold monitor was triggered, resulting in the high voltage turn off error message. The 654 status message was then issued, actually turning off the high voltage in the MCE. The path out to the NSSC-1 took slightly longer than the voltage monitoring error message. Since the hardware BSD was not triggered, the upper bound on the lamp brightness remained at less than approximately 1,000,000 counts. The Ball recommendation was to allow the MAMA2 recovery to occur nominally out of the SMS (at 088/013805Z). As agreed upon, STIS memory monitor words were set up at 086/2240Z, and the STIS event flag #3 was cleared at 087/0200Z. The NSSC-1 flight software ED report was reset at 086/1724Z, using ROP NS-5. The ground master image was updated at 088/0733Z. The SSR EDAC error counter was cleared nine times via ROP SR-1. HSTAR 6415 was written since an excessive number of EDAC counts were experienced near 088/130718Z in an unexpected block range. A re-dump of the same block range resulted in the same number of counts. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH